# Large-scale peculiar velocities in the universe

### v2

@@ 10.1.2 Redshift-space distortions
Perturbation theory expansions: These extend the Kaiser formalism into the quasi-linear regime by incorporating higher-order terms in the density and velocity fields [317, 379, 381, 382]. The Taruya-Nishimichi-Saito (TNS) model [379], for instance, supplements the Kaiser formula with correction terms and that account for higher-order coupling between density and velocity fields:
@@ 10.1.3 Contamination of CMB studies
(10.1.30) with . Also, and is the angle between the direction of the observation () and our velocity vector. The precise measurement of this dipole serves as a fundamental calibration reference for CMB experiments [422].However, recent studies have questioned whether the CMB dipole is entirely kinematic, suggesting possible contributions from large-scale structure, or even from intrinsic primordial dipoles [82].
@@ 11.1.3 Synergies with photometric surveys
The Nancy Grace Roman Space Telescope (formerly WFIRST) [561], will complement the ground-based efforts with space-based infrared imaging and spectroscopy, particularly valuable for measuring distances to galaxies in obscured regions and at higher redshifts than accessible to the current surveys.
@@ 11.3.1 Machine learning applications
Distance indicator calibration, reducing scatter in Tully-Fisher and Fundamental Plane relations [21].
Velocity field reconstruction from sparse and noisy data [98].
Identification of systematic biases in peculiar velocity catalogs [15].
Emulation of complex theoretical models for faster likelihood analysis [555].
@@ 12.3 Future prospects
Acknowledgments: This work was supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.), under the “First Call for H.F.R.I. Research Projects to support Faculty members and Researchers and the procurement of high-cost research equipment Grant” (Project Number: 789). The authors also wish to acknowledge Animesh Sah, Roya Mohayaee, Erick Pastn, Mohamed Rameez and Subir Sarkar for helpful discussions.

### v3

@@ 10.1.2 Redshift-space distortions
Perturbation theory expansions: These extend the Kaiser formalism into the quasi-linear regime by incorporating higher-order terms in the density and in the velocity fields (see [352, 412, 414, 415]). The Taruya-Nishimichi-Saito (TNS) model [412], for instance, supplements the Kaiser formula with correction terms and that account for higher-order coupling between density and velocity fields:
@@ 10.1.3 Contamination of CMB studies
(10.1.30) with . Also, and is the angle between the direction of the observation () and our velocity vector. The precise measurement of this dipole serves as a fundamental calibration reference for CMB experiments [455]. However, recent studies have questioned whether the CMB dipole is entirely kinematic, suggesting possible contributions from large-scale structure, or even from intrinsic primordial dipoles [93]. We should note that alternative approaches to the dipole effect also exist in the literature (e.g. see [456, 76]). However, the covariant tilted-frame analysis offers a unified geometric explanation for the coincidence of dipoles across different cosmological tracers.
@@ 11.1.3 Synergies with photometric surveys
The Nancy Grace Roman Space Telescope (formerly WFIRST) [594], will complement the ground-based efforts with space-based infrared imaging and spectroscopy, particularly valuable for measuring distances to galaxies in obscured regions and at higher redshifts than accessible to the current surveys. Similarly, the SPHEREx mission [595] will conduct an all-sky near-infrared spectral survey, creating a three-dimensional map of the universe by measuring redshifts for hundreds of millions of galaxies. This extensive dataset will be invaluable for mapping large-scale structure and cross-correlating with peculiar velocity surveys to reduce cosmic variance, particularly on large scales.
@@ 12.3 Future prospects
Acknowledgments: This work was supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.), under the “First Call for H.F.R.I. Research Projects to support Faculty members and Researchers and the procurement of high-cost research equipment Grant” (Project Number: 789). The authors also wish to thank Animesh Sah, Eoin Ó Colgáin, Erick Pastn, Feng Shi, Mohamed Rameez, Roya Mohayaee, Subir Sarkar, Ted Lauer and Yan-Chuan Cai for helpful discussions and comments.